OpenStudy (nurali):

A quiz consists of 10 multiple choice questions. Each question is marked either “right” or “wrong.” Each right answer gets 5 points, and 1 point is taken off for each wrong answer. In what follows, fill in the blank; if you think there is not enough information to answer the question, enter NA.

OpenStudy (nurali):

a. The correlation between the number of right answers and the number of wrong answers is _________________. b. The correlation between the score on the test and the number of right answers is _________________.

OpenStudy (anonymous):

A.? B.?

OpenStudy (anonymous):

Problem1The birthweights and lengths of newborns at a hospital have an upward sloping football shaped scatter diagram. For a baby on the 35th percentile of birthweights, the regression estimate of length a.is on the 35th percentile of lengths. b.is above the 35th percentile of lengths. c.is below the 35th percentile of lengths. d.cannot be placed relative to the 35th percentile of lengths, based on the information given.

OpenStudy (zehanz):

@Nurali: because the sum of the wrong and the right answers is 10, there is a 100% correlation between them. You can even give a formula : w = 10 - r, or r = 10 - w.

OpenStudy (zehanz):

About the second part: there is also a 100% correlation. \(Score = 5r - w\), but \(w = 10-r\), so \(Score=5r-(10-r)=6r-10\)

OpenStudy (zehanz):

I don't understand your two last responses...

OpenStudy (anonymous):

Nurali i wonder if you did this question The highest point in the scatter is the point for Illinois, at 1807. The two just below it, at 1778 and 1777, are for Minnesota and Iowa. The lowest points are at 1391 and 1385, for Maine and D.C. respectively. Pick all the options that are correct. The scatter diagram shows that... hope your promptly answer

OpenStudy (anonymous):

nez the answer is b.is above the 35th percentile of lengths.

OpenStudy (anonymous):

w=5 r =9 is wrong answer do you perhaps figure out this correlation? help please

OpenStudy (anonymous):

I pick all option and they are wrong please double check walter.

OpenStudy (anonymous):

The intercept of the regression line for estimating chick weight based on egg diameter is _______________ grams. ? The diameter of one of the eggs is 0.5 mm wider than that of another. According to the regression method, the chick that hatches from the bigger egg is estimated to be ____________ grams heavier.? Correct answer please!

OpenStudy (anonymous):

A quiz consists of 10 multiple choice questions. Each question is marked either “right” or “wrong.” Each right answer gets 5 points, and 1 point is taken off for each wrong answer. In what follows, fill in the blank; if you think there is not enough information to answer the question, enter NA. Problem 2A The correlation between the number of right answers and the number of wrong answers is _________________. Problem 2B The correlation between the score on the test and the number of right answers is _________________. Pandu wich is the correct answer?

OpenStudy (nurali):

@WalterDe The intercept of the regression line for estimating chick weight based on egg diameter is _______________ grams. ? The diameter of one of the eggs is 0.5 mm wider than that of another. According to the regression method, the chick that hatches from the bigger egg is estimated to be ____________ grams heavier.? Answer -13.159 0.416 A quiz consists of 10 multiple choice questions. Each question is marked either “right” or “wrong.” Each right answer gets 5 points, and 1 point is taken off for each wrong answer. In what follows, fill in the blank; if you think there is not enough information to answer the question, enter NA. Problem 2A The correlation between the number of right answers and the number of wrong answers is _________________. Problem 2B The correlation between the score on the test and the number of right answers is _________________. Answer -1 1

OpenStudy (anonymous):

nurali you answers are correct, how about the scatter diagram were you have to pick the best option? the same question asked by walter